OpenStudy (rebeccapink6398):

could someone please help me? Sharon has 8 cups of granola. She eats 2/3 cup each day. How long will her granola last? could someone please help me to understand and what to do? :)

OpenStudy (whpalmer4):

Let's make a different problem. Say Sharon eats 1 cup of granola each day. Can you tell me how long it will last?

OpenStudy (rebeccapink6398):

8 days and thx for helping me :D

OpenStudy (whpalmer4):

You're welcome. How about if she eats 2 cups each day?

OpenStudy (rebeccapink6398):

4 days

OpenStudy (whpalmer4):

Right. So how did you find those answers?

OpenStudy (rebeccapink6398):

umm... sorry but I don't know?? lol its just easy

OpenStudy (whpalmer4):

But what was the process?

OpenStudy (whpalmer4):

You divided the amount she had by the amount she ate each day, right?

OpenStudy (rebeccapink6398):

YES! that's a great way to put it lol

OpenStudy (whpalmer4):

Okay, well, what if she ate half a cup a day? How long would it last?

OpenStudy (rebeccapink6398):

um.. 16 days?

OpenStudy (whpalmer4):

Yep! Now, here's where it might get a bit tricky. What if she eats 2/3 of a cup each day? :-)

OpenStudy (whpalmer4):

Do you know how to divide by a fraction?

OpenStudy (rebeccapink6398):

24???????

OpenStudy (whpalmer4):

That would be the right answer, if only I had asked how long it would last if she ate 1/3 of a cup each day...

OpenStudy (whpalmer4):

But she ate twice that — 1/3 cup in the morning, and 1/3 cup at night :-)

OpenStudy (rebeccapink6398):

2/3 divied by 8= 0.0833333

OpenStudy (whpalmer4):

You have your problem backwards. She has 8 cups, and eats 2/3 of a cup each day. We can divide 8 cups into how many 2/3 cup servings?

OpenStudy (whpalmer4):

You were on the right track when you said 24, if you'd followed through with what I said... You said the granola would last for 24 days. I pointed out that that was the correct answer if she eats 1/3 of a cup each day, but in fact, she eats twice as much. If eating 1/3 of a cup each day will exhaust her supply in 24 days, if she eats twice as much each day, how long will the supply last?

OpenStudy (whpalmer4):

If you spend your money twice as fast, doesn't it only last half as long? What is half of 24 days?

OpenStudy (rebeccapink6398):

12! :D

OpenStudy (whpalmer4):

Right. Now, I think you need to learn how to divide by fractions. Good news: it's easy! Dividing by a fraction is the same as multiplying by the fraction, after you turn the fraction upside down. So, \[\frac{8}{\frac{2}{3}} = 8*\frac{3}{2} = \frac{3*8}{2} = \frac{24}{2} = 12\]

OpenStudy (whpalmer4):

Or to use another example that we touched: \[\frac{8}{\frac{1}{2}} = 8*\frac{2}{1} = \frac{8*2}{1} =\frac{16}{1} = 16\]

OpenStudy (anonymous):

yup its 12

OpenStudy (anonymous):

awesome